# RapVerse: Coherent Vocals and Whole-Body Motions Generations from Text

### v1

@@ 4.2 Motion VQ-VAE Tokenizer
$Q(z^{\mathcal{M}};\mathcal{C}^{\mathcal{M}})=\mathop{\arg\min}_{k}\|z-c_{k}^{\mathcal{M}}\|_{2}.$ (3)
@@ 4.3 Vocal2unit Audio Tokenizer
Overall, we leverage the self-supervised framework polyak2021speech in speech resynthesis domain to learn vocal representations from the audio sequences. Specifically, we train a Vocal2unit audio tokenizer to build a discrete tokenized representation for the human singing voice. The vocal tokenizer consists of three encoders and a vocoder. The encoders include three different parts: (1) the semantic encoder; (2) the F0 encoder; and (3) the singer encoder. We will introduce each component of the model separately.
@@
Semantic Encoder. In order to extract the semantic information of the audio into discrete tokens, we use a pre-trained Hubert encoder hsu2021hubert as the main component of our semantic encoder network $E^{h}$. Given a raw audio sequence $v$, the output of the Hubert encoder is continuous representation. Then we apply a K-means algorithm on the representations to finally generate discrete units $z^{\mathcal{S}}=\{z^{\mathcal{S}}_{i}\}_{i=1}^{L_{s}}$ where $z^{\mathcal{S}}_{i}\in\{1,2,\cdots,K\}$. $L_{s}$ is the length of hubert token sequence, and $K$ represents the number of clusters in K-means algorithm.
@@
F0 Encoder. We use the YAAPT algorithm kasi2002yet to extract the F0 from the input audio signal $v$, which is then fed into an encoder $E_{p}$ to generated low-frequency F0 pitch representation. The encoder $E^{p}$ is trained using a VQ-VAE framework with Exponential Moving Average updates similar to dhariwal2020jukebox , maintaining a learnable codebook $\mathcal{C}^{p}=\{c_{k}^{p}\}_{k=1}^{d_{c}}$, where $d_{c}$ is the number of codes in the codebook. The raw F0 sequence is first preprocessed according to the singer, converted into a sequence of latent vectors, then mapped to its nearest neighbors in the codebook $c^{p}$, which could be represented using integer number $k\in\{1,2,\cdots,d_{c}\}$ as its index in the codebook. The index sequence will be further formatted as the token <pitch_k> to produce pitch token sequence $z^{\mathcal{P}}=\{z^{\mathcal{P}}_{i}\}_{i=1}^{L_{p}}$, where $L_{p}$ is the length of the pitch token sequence. Finally, the decoder will reconstruct the F0 representation leveraging the pitch token sequence.
@@
Singer Encoder. As the last component of our encoders, the singer encoder $E_{si}$ from heigold2016end ; polyak2021speech is used to extract the Mel-spectrogram from the raw audio sequence and to output a singer representation $z^{\mathcal{I}}\in\mathbb{R}^{256}$. The singer embedding only depends on the singer and is trained globally.
@@
Vocoder. In order to decode the vocal signal from the discrete tokens, we adopt a modified version of the HiFi-GAN neural vocoder similar to the one in polyak2021speech , consisting of a generator and multiple discriminators. The generator takes the encoded discrete representation $z^{\mathcal{S}}$ and $z^{\mathcal{P}}$, and the singer embedding vector $z^{\mathcal{I}}$ as inputs. After a lookup table and a series of blocks composed of transposed convolution and a residual block with dilated layers, it finally outputs waveform audio. The discriminators include multiple sub-discriminators, analyzing on different scales and different periodic structures.
@@ 4.4 General Auto-regressive Modeling
Model Architectures. The whole model of our general modeling pipeline consists of several pre-trained vocal and motion tokenizers and a large Text-Motion-Audio foundation model. The vocal and motion tokenizers are based on VQ-VAE architectures, enabling us to represent these two modalities as discrete tokens. We also adopt a T5-Tokenizer raffel2020exploring to convert lyric text into tokens, so that all three modalities $\mathcal{X}\in\{\mathcal{L},\mathcal{V},\mathcal{M}\}$ are united in the token space $\mathcal{T}^{\mathcal{X}}=\{t_{k}^{\mathcal{X}}\}_{k=1}^{\mathcal{N}^{\mathcal{X}}}$, where $\mathcal{N}^{\mathcal{X}}$ is the token sequence length for each modality. The large foundation model, a temporal transformer based on the decoder-only architecture, performs the next-token-prediction task based on textual features encoded by the T5-Encoder.
@@
$-\sum\limits_{i=1}^{\mathcal{N^{\mathcal{X}}}}\log p_{\delta}(t^{\mathcal{X}}_{i};h^{\mathcal{L}}).$ (5)
@@
After optimizing via this training objective, our model learns to predict the next token, which can be decoded into different modality features. This process is similar to text word generation in language models, while the “word” in our method such as <face_0123>, does not have explicit semantic information, but can be decoded into continuous modality features.
@@
Multimodality Tokens Organization. Since we have designed a coherent vocal and motion generation pipeline with multiple token types in our method, including hubert and pitch tokens in rap vocals, and face, body, hand in whole-body motions, the organization method of all these tokens in one sentence is important for training. Similar to previous works lu2023humantomato ; jiang2024motiongpt , we adopt an interleaved style within each modality. Specifically, with the input whole-body motion tokens $\mathcal{T}^{\mathcal{M}_{s}}=\{t^{\mathcal{M}_{s}}_{k}\}_{k=1}^{\mathcal{N}^{\mathcal{M}_{s}}}$, where $s=f,b,h$ which denotes face, body and hand respectively, and $\mathcal{N}^{\mathcal{M}_{s}}$ denotes the token length for each part. Therefore the motion token sequence can be formulated as $\mathcal{T}^{\mathcal{M}}=\{t^{\mathcal{M}_{f}}_{1},t^{\mathcal{M}_{b}}_{1},t^{\mathcal{M}_{h}}_{1},t^{\mathcal{M}_{f}}_{2},...\}$. Similarly, the vocal tokens consist of the hubert tokens $t^{\mathcal{V}_{h}}$ and pitch tokens $t^{\mathcal{V}_{p}}$, which can be formulated as $\mathcal{T}^{\mathcal{V}}=\{t^{\mathcal{V}_{h}}_{1},t^{\mathcal{V}_{p}}_{1},t^{\mathcal{V}_{h}}_{2},t^{\mathcal{V}_{p}}_{2},...\}$. This interleaved way of organization allows us to align the timesteps of each part. When performing coherent vocal and motion generation, we mix these two modality sequences sequentially with modality-wise start token as $\mathcal{T}=\{t_{start}^{V},\mathcal{T}^{V},t_{start}^{M},\mathcal{T}^{M}\}$. Note that we put vocal sequences before motion sequences because 1) the vocals are more related to the lyrics, and 2) motion tokens have a relationship with previously generated vocals, e.g. face motions with lips movement are directly related to vocals. Joint training of mixed modalities allows the transformer to capture the temporal dependencies and dynamics inherent in vocal and motion data.
@@
Inference and Decoupling. In the inference stage, we use different start tokens to specify which modality to generate. The textual input is encoded as features to guide token inference. We also adopt a top-k algorithm to control the diversity of the generated content by adjusting the temperature, as generating vocals and motions based on lyrics is a creation process with multiple possible answers. After token prediction, a decoupling algorithm is used to process output tokens to make sure tokens from different modalities are separated and temporally aligned. These discrete tokens will be further decoded into text-aligned vocals and motions.
@@ 5.1 Experimental Setup
Metrics. To evaluate the generation quality of singing vocals, we utilize the Mean Opinion Score (MOS) to gauge the naturalness of the synthesized vocal. For motion synthesis, we evaluate the generation quality of the body hand gestures and the realism of the face, respectively. Specifically, for gesture generation, we use Frechet Inception Distance (FID) based on a feature extractor from guo2022generating to evaluate the distance of feature distributions between the generated and real motions, and Diversity (DIV) metric to assess the motions diversity. For face generation, we compare the vertex MSE xing2023codetalker and the vertex L1 difference LVD yi2023generating . Finally, we adopt Beat Constancy (BC) li2021ai to measure the synchrony of generated motion and singing vocals.
Baselines. We compare the vocal generation quality with the state-of-the-art vocal generation method DiffSinger liu2022diffsinger . And we also adapt the text-to-speech model FastSpeech2 ren2020fastspeech for vocal generation. For motion generation, we compare our method with both text-to-motion methods and audio-to-motion methods. For text-to-motion methods, since there is no existing open-sourced work for text to whole-body motion generation, we compare with transformer-based T2M-GPT zhang2023t2m and MLD chen2023executing for body generation. For the audio-to-motion generation, we compare with Habibie et al. habibie2021learning and the SOTA model Talkshow yi2023generating . We report all the results on RapVerse with an 85%/7.5%/7.5% train/val/test split.
@@
T2M-GPT zhang2023t2m 18.42 11.75 - - -
MLD chen2023executing 24.52 12.15 - - -
@@
Habibie et al. habibie2021learning 32.14 10.08 0.476 2.11 9.54
Talkshow yi2023generating 21.10 13.14 0.487 1.93 9.20
Cascaded Result 23.42 12.87 0.479 2.05 9.38
@@
FastSpeech2 ren2020fastspeech 3.41 ± 0.18
DiffSinger liu2022diffsinger 3.72 ± 0.12
@@ 5.2 Main Results Analysis
Furthermore, the cornerstone of our approach lies in the simultaneous generation of vocals and motion, aiming to achieve temporal alignment between these modalities. This objective is substantiated by our competitive results on the BC metric, which assesses the synchrony between singing vocals and corresponding motions, underscoring our success in closely synchronizing the generation of these two modalities. For the cascaded system, we integrate the text-to-vocal model DiffSinger with the audio-to-motion model Talkshow. Compared with the cascaded system, our joint-generation pipeline demonstrates superior outcomes while also reducing computational demands during both training and inference phases. In the cascaded architectures, errors tend to accumulate through each stage. Specifically, if the text-to-vocal module produces unclear vocals, it subsequently hampers the audio-to-motion model’s ability to generate accurate facial expressions that align with the vocal content.
Evaluations on vocal generations. We have also compared our method against other state-of-the-art text-to-vocal generation baselines in Table. 2 (b). While our unified model is trained to simultaneously generate vocals and motion, a task considerably more complex than generating vocals alone, its vocal generation component still manages to achieve results comparable to those systems designed solely for vocal generations.
@@
GT 0 12.53 0.499 0 0
Pretrained LLM 48.25 10.05 0.462 3.28 12.15
Single motion token 19.15 11.29 0.477 2.18 10.12
Ours 18.65 12.19 0.481 2.03 7.23
@@ 5.3 Ablation Study
We present the outcomes of our ablation study in Table. 3. Initially, we explored the integration of a pre-trained large language model raffel2020exploring for multi-modality generation, akin to the approach in  jiang2024motiongpt . However, the efficacy of utilizing pre-trained language models significantly lags behind our tailored design, underscoring that pre-training primarily on linguistic tokens does not facilitate effective prediction across multiple modalities like vocal and motion. Additionally, we study the impact of our compositional VQ-VAEs on motion generation. In contrast, a baseline employing a single VQVAE for the joint quantization of facial, body, and hand movements was implemented. This approach led to a noticeable degradation in performance, particularly marked by a -2.89 decrease in LVD. This decline can be attributed to the preponderance of facial movements in a singer’s performance. Using a single VQ-VAE model for full-body dynamics compromises the detailed representation of facial expressions, which are crucial for realistic and coherent motion synthesis.
@@ 6 Conclusion
Limitations and Future Works. One limitation of RapVerse is that it is currently constrained to the rap music generation, excluding other music genres. However, we want to emphasize that our method is a general framework that could be flexibly utilized in other scenarios for joint audio and motion generation if provided with the datasets. Another important future direction resides in multi-performer audio and motion generation, which could be utilized in contemporary music performances such as virtual live bands.
Appendix A Appendix
The supplementary material is organized as follows: Sec. A.1 provides a website demo to show additional qualitative results; Sec. A.2 presents additional details of the network architectures; Sec. A.3 introduces evaluation metrics; Sec. A.4 presents additional ablation studies; Sec. A.5 shows additional qualitative results; Sec. A.6 discusses broader societal impacts of the work.
A.1 Website Demo
In order to provide more vivid and clear qualitative results, we make a supplemental website demo to demonstrate the generation quality of our proposed system. We encourage readers to view the results at https://vis-www.cs.umass.edu/RapVerse/.
A.2 Implementation Details
Motion Tokenizer. We train three separate Vector Quantized Variational Autoencoders (VQ-VAE) for face, body and hand, respectively. We adopt the same VQ-VAE architecture based on guo2022tm2t ; zhang2023t2m ; jiang2024motiongpt . For loss functions optimizing the motion tokenizers, we use the L1 smooth reconstruction loss, the embedding loss, and the commitment loss. The commitment loss weight is set to 0.02. In line with guo2022tm2t ; jiang2024motiongpt , strategies such as the exponential moving average and the codebook reset technique razavi2019generating are implemented to optimize codebook efficacy throughout the training process. We take 512 for the codebook size and set the dimension of each code to 512. We set the temporal down-sampling rate to 4. We train the VQ-VAEs with a batch size of 256 and a sequence window length of 72. We adopt Adam with $\beta_{1}=0.9$, $\beta_{2}=0.99$, and a learning rate of 0.0002 as the optimizer.
Vocal Tokenizer. For the semantic encoder, we adopt a BASE 12 transformer of HuBERT hsu2021hubert pre-trained on the 969-hour LibriSpeech corpus panayotov2015librispeech . Following lakhotia2021generative ; polyak2021speech , we derive the feature activations from its sixth layer. This process allows the HuBERT model to transform input audio into a 768-dimensional vector space. Subsequently, we employ the k-means algorithm with 500 centroids to get quantized discrete content codes. For the F0 encoder, a VQ-VAE framework is utilized to discretize the F0 signal into quantized F0 tokens. We adopt the Exponential Moving Average updates during training the VQ-VAE following dhariwal2020jukebox ; polyak2021speech . We set the codebook size of the VQ-VAE to 20 entries. Moreover, as the original work directly normalizes the extracted F0 values for each singer respectively, we don’t explicitly use the singer’s statistics but adopt a windowed convolutional layer with both the audio input (sliced into the window size) and singer’s embedding as input. Finally, we adopt a similar architecture as heigold2016end for the singer encoder.
General Auto-regressive Model. The auto-regressive model consists of a T5 Embedder and a Foundation Model. We use a T5-Large Encoder as our Embedder, with 24 layers and 16 heads. The Embedder is freezed during foundation model training. The foundation model is based on the Decoder-only transformer architecture, which has 12 layers and 8 heads. We use Adam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.99$, and a learning rate of 0.0002. We do not use dropout in our training. Our training batch size is 384 for 100 epochs.
A.3 Evaluation Metrics
To evaluate the motion generation quality, we utilize the following metrics:
  1. 1.
    Frechet Inception Distance (FID): This metric measures the distribution discrepancy between the ground truth and generated motions of body and hand gestures. Specifically, we train an autoencoder based on guo2022tm2t as the motion feature extractor.
  2. 2.
    Diversity (DIV): DIV evaluates the diversity of generated motions, where we calculate the variance from the extracted motion features.
  3. 3.
    Vertex MSE: Following xing2023codetalker , we compute the mean L2 error of lip vertices between generated face motions and ground truth.
  4. 4.
    Landmark Velocity Difference (LVD): Introduced by yi2023generating , LVD calculates the velocity difference of generated facial landmarks and ground truth.
  5. 5.
    Beat Constancy (BC) li2021ai : BC evaluates the synchrony of generated motions and singing vocals by calculating the similarity between the rhythm of gestures and audio beat. Specifically, we extract audio beats using librosa mcfee2015librosa , and we compute the kinematic beats as the local minima of the velocity of joints. Then the alignment score is derived from the mean proximity of motion beats to the nearest audio beat.
For the evaluation of singing vocal generation quality, the Mean Opinion Score (MOS) is employed. It reflects the perceived naturalness of the synthesized vocal tones, with human evaluators rating each sample on a scale from 1 to 5, thereby offering a subjective measure of vocal synthesis fidelity.
A.4 Additional Ablation Studies
Ablation on Motion Tokenizer. We study different designs of our motion tokenizer by comparing the reconstruction results. Specifically, we explore VQ-VAEs with different codebook sizes, and study the effect of using a single VQ-VAE to model full-body motions instead of multiple VQ-VAEs for different body parts. As is demonstrated in Table. 4, we find that using separate VQ-VAEs for face, body and hands has lower reconstruction error. And we select a codebook size of 512 for our final model.
[…]
K = 256 63.4 39.1 9.14
K = 512 60.9 35.4 8.90
K = 1024 62.2 36.8 9.08
Single 65.2 38.8 9.32
Ours 60.9 35.4 8.90
Table 4: Evaluation of our motion tokenizer. We follow lu2023humantomato to evaluate the motion reconstruction errors of our motion VQVAE model $V_{m}$. MPJPE and PAMPJPE are measured in millimeters. ACCL indicates acceleration error.
Ablation on Vocal Tokenizer. We also study different designs for our audio tokenizer by comparing the reconstruction results. Specifically, we explore different codebook sizes for the semantic encoder by changing the K-Means number. We also compare the effect with our singer embedding in F0 value postprocessing. We use the following metrics to measure the reconstruction quality of the vocal tokenizer:
  1. 1.
    Character Error Rate (CER): We use Whisper radford2022whisper to transcribe the ground truth and synthesized audios, and then take the corresponding ground truth lyrics as the reference to calculate the CER of the synthesized audios.
  2. 2.
    Gross Pitch Error (GPE): The percentage of pitch estimates that have a deviation of more than 20% around the true pitch. Only frames considered pitched by both the ground truth and the synthesized audio are considered.
  3. 3.
    Voicing Decision Error (VDE) NAKATANI2008203 : The portion of frames with voicing decision error, that is, the results using ground truth and synthesized audio to determine whether the frame is voiced are different.
[…]
GT 41.88 - -
K = 100 76.39 2.64 11.98
K = 500 (Ours) 69.21 2.29 8.84
K = 2000 67.46 2.54 9.08
Original F0 post-process 68.93 2.71 9.19
Table 5: Evaluation of our unit2wav model. We follow polyak2021speech to evaluate the speech resynthesis errors of our unit2wav model. WER, GPE, and VDE, expressed as percentages, indicate the character error rate, the grand pitch error and the voicing decision error.
The analysis of the results indicates that even the original audio exhibits a high CER, which could be attributed to the rapid speech rate associated with rapping. In some instances, the lyrics may not be distinctly recognizable even by human listeners. Upon comparing different codebook sizes, it is observed that they achieve comparable GPE values. This similarity in GPE is expected since the same F0 model is employed across all codebook sizes. The CER, which serves as a direct measure of the semantic information preserved in the code, suggests that larger codebooks tend to retain more semantic information. However, the difference in CER between codebook sizes of $K=500$ and $K=2000$ is minimal. Given that $K=500$ demonstrates better GPE and VDE, we select $K=500$.
Additionally, we ablate a design without the singer embedding in the F0 preprocessing, instead normalizing the F0 values for each singer. It shows that this approach resulted in significantly inferior performance, particularly in pitch prediction, compared to the modified version that includes the singer embedding.
A.5 Additional Qualitative Results
We show additional qualitative results in Fig. 4. Our model adeptly generates comprehensive whole-body motions that embody the essence of the input lyrics. These include authentic gesture movements that resonate with the song’s rhythm and synchronized lip motions that articulate the lyrics.
Figure 4: Additional qualitative results. Our method can generate diverse whole-body motions from input lyrics.
A.6 Broader Impacts
This research contributes to advancements in generating synchronized vocals and human motion from textual lyrics, aiming to enhance virtual agents’ ability to provide immersive and interactive experiences in digital media. The potential positive impact of this work lies in its ability to create more lifelike and engaging virtual performances, such as in virtual concerts and gaming, where characters can perform and react in ways that are deeply resonant with human expressions. This could significantly enhance user engagement in virtual reality settings, and provide innovative solutions in entertainment industries.
However, this capability carries inherent risks of misuse. The technology’s ability to generate realistic human-like actions and singing vocals from mere text raises concerns about its potential to create misleading or deceptive content. For example, this could be exploited to produce fake videos or deepfakes where individuals appear to sing and perform that never actually occurred, which could be used to spread misinformation or harm reputations. Recognizing these risks, it is crucial to advocate for ethical guidelines and robust frameworks to ensure the responsible use of such technologies.

### v3

@@ 4.2 Motion VQ-VAE Tokenizer
$Q(z^{\mathcal{M}};\mathcal{C}^{\mathcal{M}})=\mathop{\arg\min}_{k}\|z^{\mathcal{M}}-c_{k}^{\mathcal{M}}\|_{2}.$ (3)
@@ 4.3 Vocal2unit Audio Tokenizer
Overall, we leverage the self-supervised framework [polyak2021speech] in speech resynthesis domain to learn vocal representations from the audio sequences. Specifically, we train a Vocal2unit audio tokenizer to build a discrete tokenized representation for the human singing voice. The vocal tokenizer consists of three encoders and a vocoder. The encoders include three different parts: (1) the semantic encoder; (2) the F0 encoder; and (3) the singer encoder. We will introduce each component of the model separately.
@@
Semantic Encoder. In order to extract the semantic information of the audio into discrete tokens, we use a pre-trained Hubert encoder [hsu2021hubert] as the main component of our semantic encoder network $E^{h}$. Given a raw audio sequence $v$, the output of the Hubert encoder is a continuous representation. Then we apply a K-means algorithm on the representations to finally generate discrete units $z^{\mathcal{S}}=\{z^{\mathcal{S}}_{i}\}_{i=1}^{L_{s}}$ where $z^{\mathcal{S}}_{i}\in\{1,2,\cdots,K\}$. $L_{s}$ is the length of hubert token sequence, and $K$ represents the number of clusters in K-means algorithm.
@@
F0 Encoder. We use the YAAPT algorithm [kasi2002yet] to extract the F0 from the input audio signal $v$, which is then fed into an encoder $E_{p}$ to generated low-frequency F0 pitch representation. The encoder $E^{p}$ is trained using a VQ-VAE framework with Exponential Moving Average updates similar to [dhariwal2020jukebox], maintaining a learnable codebook $\mathcal{C}^{p}=\{c_{k}^{p}\}_{k=1}^{d_{c}}$, where $d_{c}$ is the number of codes in the codebook. The raw F0 sequence is first preprocessed according to the singer, converted into a sequence of latent vectors, then mapped to its nearest neighbors in the codebook $c^{p}$, which could be represented using integer number $k\in\{1,2,\cdots,d_{c}\}$ as its index in the codebook. The index sequence will be further formatted as the token ¡pitch_k¿ to produce pitch token sequence $z^{\mathcal{P}}=\{z^{\mathcal{P}}_{i}\}_{i=1}^{L_{p}}$, where $L_{p}$ is the length of the pitch token sequence. Finally, the decoder will reconstruct the F0 representation leveraging the pitch token sequence.
@@
Singer Encoder. As the last component of our encoders, the singer encoder $E_{si}$ from [heigold2016end, polyak2021speech] is used to extract the Mel-spectrogram from the raw audio sequence and to output a singer representation $z^{\mathcal{I}}\in\mathbb{R}^{256}$. The singer embedding only depends on the singer and is trained globally.
@@
Vocoder. In order to decode the vocal signal from the discrete tokens, we adopt a modified version of the HiFi-GAN neural vocoder similar to the one in [polyak2021speech], consisting of a generator and multiple discriminators. The generator takes the encoded discrete representation $z^{\mathcal{S}}$ and $z^{\mathcal{P}}$, and the singer embedding vector $z^{\mathcal{I}}$ as inputs. After a lookup table and a series of blocks composed of transposed convolution and a residual block with dilated layers, it finally outputs waveform audio. The discriminators include multiple sub-discriminators, analyzing on different scales and different periodic structures.
@@ 4.4 General Auto-regressive Modeling
Model Architectures. The whole model of our general modeling pipeline consists of several pre-trained vocal and motion tokenizers and a large Text-Motion-Audio foundation model. The vocal and motion tokenizers are based on VQ-VAE architectures, enabling us to represent these two modalities as discrete tokens. We also adopt a T5-Tokenizer [raffel2020exploring] to convert lyric text into tokens, so that all three modalities $\mathcal{X}\in\{\mathcal{L},\mathcal{V},\mathcal{M}\}$ are united in the token space $\mathcal{T}^{\mathcal{X}}=\{t_{k}^{\mathcal{X}}\}_{k=1}^{\mathcal{N}^{\mathcal{X}}}$, where $\mathcal{N}^{\mathcal{X}}$ is the token sequence length for each modality. The large foundation model, a temporal transformer based on the decoder-only architecture, performs the next-token-prediction task based on textual features encoded by the T5-Encoder.
@@
$\sum\limits_{i=1}^{\mathcal{N^{\mathcal{X}}}}\log p_{\delta}(t^{\mathcal{X}}_{i};h^{\mathcal{L}}).$ (5)
@@
After optimizing via this training objective, our model learns to predict the next token, which can be decoded into different modality features. This process is similar to text word generation in language models, while the “word” in our method such as ¡face_0123¿, does not have explicit semantic information, but can be decoded into continuous modality features.
@@
Multimodality Tokens Organization. Since we have designed a coherent vocal and motion generation pipeline with multiple token types in our method, including hubert and pitch tokens in rap vocals, and face, body, hand in whole-body motions, the organization method of all these tokens in one sentence is important for training. Similar to previous works [lu2023humantomato, jiang2024motiongpt], we adopt an interleaved style within each modality. Specifically, with the input whole-body motion tokens $\mathcal{T}^{\mathcal{M}_{s}}=\{t^{\mathcal{M}_{s}}_{k}\}_{k=1}^{\mathcal{N}^{\mathcal{M}_{s}}}$, where $s=f,b,h$ which denotes face, body and hand respectively, and $\mathcal{N}^{\mathcal{M}_{s}}$ denotes the token length for each part. Therefore the motion token sequence can be formulated as $\mathcal{T}^{\mathcal{M}}=\{t^{\mathcal{M}_{f}}_{1},t^{\mathcal{M}_{b}}_{1},t^{\mathcal{M}_{h}}_{1},t^{\mathcal{M}_{f}}_{2},$...\}. Similarly, the vocal tokens consist of the hubert tokens $t^{\mathcal{V}_{h}}$ and pitch tokens $t^{\mathcal{V}_{p}}$, which can be formulated as $\mathcal{T}^{\mathcal{V}}=\{t^{\mathcal{V}_{h}}_{1},t^{\mathcal{V}_{p}}_{1},t^{\mathcal{V}_{h}}_{2},t^{\mathcal{V}_{p}}_{2},$...\}. This interleaved way of organization allows us to align the timesteps of each part. When performing coherent vocal and motion generation, we mix these two modality sequences sequentially with modality-wise start token as $\mathcal{T}=\{t_{start}^{V},\mathcal{T}^{V},t_{start}^{M},\mathcal{T}^{M}\}$. Note that we put vocal sequences before motion sequences because 1) the vocals are more related to the lyrics, and 2) motion tokens have a relationship with previously generated vocals, e.g. face motions with lips movement are directly related to vocals. Joint training of mixed modalities allows the transformer to capture the temporal dependencies and dynamics inherent in vocal and motion data.
@@
Inference and Decoupling. In the inference stage, we use different start tokens to specify which modality to generate. The textual input is encoded as features to guide token inference. We adopt a top-k algorithm to control the diversity of the generated content by adjusting the temperature, as generating vocals and motions based on lyrics is a creation process with multiple possible answers. After token prediction, a decoupling algorithm is used to process output tokens to make sure tokens from different modalities are separated and temporally aligned. These discrete tokens will be further decoded into text-aligned vocals and motions.
@@ 5.1 Experimental Setup
Metrics. To evaluate the generation quality of singing vocals, we utilize the Mean Opinion Score (MOS) to gauge the naturalness of the synthesized vocal. For motion synthesis, we evaluate the generation quality of the body hand gestures and the realism of the face, respectively. Specifically, for gesture generation, we use Frechet Inception Distance (FID) based on a feature extractor from [guo2022generating] to evaluate the distance of feature distributions between the generated and real motions, and Diversity (DIV) metric to assess the motions diversity. For face generation, we compare the vertex MSE [xing2023codetalker] and the vertex L1 difference LVD [yi2023generating]. Finally, we adopt Beat Constancy (BC) [li2021ai] to measure the synchrony of generated motion and singing vocals.
Baselines. We compare the vocal generation quality with the state-of-the-art vocal generation method DiffSinger [liu2022diffsinger]. And we also adapt the text-to-speech model FastSpeech2 [ren2020fastspeech] for vocal generation. For motion generation, we compare our method with both text-to-motion methods and audio-to-motion methods. For text-to-motion methods, since there is no existing open-sourced work for text to whole-body motion generation, we compare with transformer-based T2M-GPT [zhang2023t2m] and MLD [chen2023executing] for body generation. For the audio-to-motion generation, we compare with Habibie et al. [habibie2021learning], Talkshow [yi2023generating], Motioncraft [bian2024adding] and EMAGE [liu2023emage]. We report all the results on RapVerse with an 85%/7.5%/7.5% train/val/test split.
@@
T2M-GPT [zhang2023t2m] 23.45 11.75 - - -
MLD [chen2023executing] 26.34 12.15 - - -
@@
Habibie et al. [habibie2021learning] 32.14 10.08 0.476 2.13 9.54
Talkshow [yi2023generating] 18.23 13.14 0.482 2.05 9.20
EMAGE [liu2023emage] 21.18 12.65 0.488 1.96 8.45
MotionCraft [bian2024adding] 17.75 13.76 0.482 2.06 9.23
Cascaded Result 23.42 12.87 0.479 2.09 9.38
@@
FastSpeech2 [ren2020fastspeech] 3.41 ± 0.18
DiffSinger [liu2022diffsinger] 3.72 ± 0.12
@@ 5.2 Main Results Analysis
GT 0 14.53 0.499 0 0
Pretrained LLM 48.25 12.65 0.462 3.28 12.15
Single motion token 19.15 12.75 0.477 2.18 10.12
Ours 17.58 14.08 0.485 2.03 7.23
@@
Figure 4: User study on generated motion.
Furthermore, the cornerstone of our approach lies in the simultaneous generation of vocals and motion, aiming to achieve temporal alignment between them. This objective is substantiated by our competitive results on the BC metric, which assesses the synchrony between singing vocals and corresponding motions, underscoring our success in closely synchronizing the generation of these two modalities. For the cascaded system, we integrate the text-to-vocal model DiffSinger with the audio-to-motion model Talkshow. Compared with the cascaded system, our joint-generation pipeline demonstrates superior outcomes while also reducing computational demands during both training and inference phases. In the cascaded architectures, errors tend to accumulate through each stage. Specifically, if the text-to-vocal module produces unclear vocals, it subsequently hampers the audio-to-motion model’s ability to generate accurate facial expressions that align with the vocal content.
To better evaluate the generation quality, we conducted an additional user study (see Fig. 4) using ten 10-second-long videos from the RapVerse dataset, sampled from the test set. We recruited 20 participants to evaluate four aspects: the holistic realism of human motion, expression-vocal synchronism, gesture-vocal synchronism, and holistic diversity, using real data as a reference. Participants were asked to sort shuffled videos generated by different methods based on these criteria. As shown in the results, our model achieved the highest scores in human motion realism, gesture-vocal synchronism, and holistic diversity when compared with T2MGPT (text-to-motion) and Talkshow (audio-to-motion).
Evaluations on vocal generations. We compare our method against other state-of-the-art text-to-vocal generation baselines in Table. 2 (b). While our unified model is trained to simultaneously generate vocals and motion, a task considerably more complex than generating vocals alone, its vocal generation component still manages to achieve results comparable to systems designed solely for vocal generations.
@@ 5.3 Ablation Study
We present the outcomes of our ablation study in Table. 3. Initially, we explored the integration of a pre-trained large language model [raffel2020exploring] for multi-modality generation, akin to the approach in  [jiang2024motiongpt]. However, the efficacy of utilizing pre-trained language models significantly lags behind our tailored design, underscoring that pre-training primarily on linguistic tokens does not facilitate effective prediction across multiple modalities like vocal and motion. Additionally, we study the impact of our compositional VQ-VAEs on motion generation. In contrast, a baseline employing a single VQVAE for the joint quantization of facial, body, and hand movements was implemented. This approach led to a noticeable degradation in performance, particularly marked by a -2.89 decrease in LVD. This decline can be attributed to the preponderance of facial movements in a singer’s performance. Using a single VQ-VAE model for full-body dynamics compromises the detailed representation of facial expressions, which are crucial for realistic and coherent motion synthesis.